# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ Abstract
Keywords: Topic Modeling, Latent Dirichlet Allocation, Total Variation Penalty, Graph Regularization, Graph Cross Validation;
@@ 1 Introduction
Originally developed to reduce document representations to low-dimensional latent semantic spaces, topic modeling has been successfully deployed for the analysis of count data in a number of applications, ranging from image processing (Tu et al., 2016; Zheng et al., 2015) and image annotation  (Feng and Lapata, 2010; Shao et al., 2009), to biochemistry (Reder et al., 2021), genetics (Dey et al., 2017; Kho et al., 2017; Liu et al., 2016; Yang et al., 2019) or microbiome studies (Sankaran and Holmes, 2019; Symul et al., 2023).
@@
    Analyzing tumor microenvironments: In this context, slices of tumor samples are partitioned into smaller regions known as “tumor microenvironments,” where the frequency of specific immune cell types is recorded (Chen et al., 2020). Here, documents correspond to tumor microenvironments, and words to cell types. The objective is to identify communities of co-abundant cells—the “topics”, taken here as a proxy for patterns of tumor-immunecell interactions— to evaluate whether some can be associated with specific outcomes such as survival rates. In this setting, a reasonable hypothesis consists in assuming that neighboring microenvironments share similar topic proportions. Since these microenvironments are inherently small, leveraging the spatial smoothness of the mixture matrix $W$ can significantly improve inference (Chen et al., 2020). We develop this example in further detail in Section 4.1.
@@
    Microbiome studies: Topic models have also been proven to be extremely useful in microbiome analysis (Sankaran and Holmes, 2019; Symul et al., 2023). In this setting, the data consists of a microbiome count matrix recording the amount of different types of bacteria found in each sample. In this case, microbiome samples are identified to documents, with bacteria playing the roles of the vocabulary, and the goal becomes to identify communities of co-abundant bacteria (Sankaran and Holmes, 2019). When additional covariate information is available (such as age, gender, and other demographic details), we can expect similar samples (documents) to exhibit similar community compositions (topic proportions).
@@
    The analysis of short documents, such as a collection of scientific abstracts or recipes: In this case, while recipes might be short, information on the origin of the recipe can help determine the topics and mixture matrix more accurately by leveraging the assumption that recipes wof neighboring countries will typically share similar topic proportions. We elaborate on this example in greater detail in Section 4.2.
Prior works. To the best of our knowledge, previous attempts to incorporate metadata to improve topic estimation have focused on Bayesian approaches to topic modeling, usually extending the latent Dirichlet allocation (LDA) model of Blei et al. (2001). We defer the detailed description of these methods to Section 5, and only briefly summarize their main characteristics here. By and large, these methods typically incorporate the metadata —often in the form of a covariate matrix—within a prior distribution (Blei and Lafferty, 2006a, b; Roberts et al., 2014; Mcauliffe and Blei, 2007). However, these models (a) are difficult to adapt to different types of covariates or information encoded as a graph, and (b) typically lack theoretical guarantees. Recent work by Chen et al. (2020) proposes extending LDA to analyze documents with known similarities by smoothing the topic proportion hyperparameters along the edges of the graph. However, this method does not empirically yield spatially smooth structures (see Sections  3.4 and  4), and significantly increases the algorithm’s running time.
In the frequentist realm, probabilistic latent semantic indexing (pLSI) has benefited from renewed interest over the past five years. Like LDA, pLSI is built around the data generation mechanism explicated in (1), effectively modeling documents as bags of words. Unlike LDA, pLSI assumes that the matrices $A$ and $W$ are fixed parameters, instead of random variables. In particular, new work by Ke and Wang (2017) and Klopp et al. (2021) suggest procedures to reliably estimate the topic matrix $A$ and mixture matrix $W$, respectively, and rigorously characterize the consistency and efficiency of their estimators through high-probability error bounds. Although recent work has begun investigating the use of structures to improve estimation in pLSI-based topic models, most approaches have limited this to considering various versions of sparsity (Bing et al., 2020; Wu et al., 2023) or weak sparsity (Tran et al., 2023) for the topic matrix $A$. To the best of our knowledge, however, no pLSI approach has yet been proposed that effectively leverages similarities between documents nor characterizes the consistency of these estimators.
@@
We propose a scalable algorithm based on a graph-aligned singular decomposition of the empirical frequency matrix $X$ to provide estimates of $W$ and $A$ (Section 2). Additionally, we develop a new cross-validation procedure for our graph-based penalty that allows us to choose the optimal regularization parameter adaptively (Section 2.3).
@@
We prove the benefits of the graph alignment procedure by deriving high probability upper bounds for both $W$ and $A$ (Section 3), which we verify through extensive simulations (Section 3.4).
@@
We showcase the advantage of our method over LDA-based methods and non-structured pLSI techniques on three real-world datasets: two spatial transcriptomics examples (Section 4.1) and a recipe dataset (Section 4.2).
@@
Throughout this paper, we use the following notations. For any $t\in\mathbb{Z}_{+}$, $[t]$ denotes the set $\{1,2,...,t\}$. For any $a,b\in\mathbb{R}$, we write $a\vee b=\max(a,b)$ and ${a\wedge b=\min(a,b)}$. We use $\mathbf{1}_{d}\in\mathbb{R}^{d}$ to denote the vector with all entries equal to 1 and $\mathbf{e}_{k}\in\mathbb{R}^{d}$ to denote the vector with $k^{th}$ element set to 1 and 0 otherwise. For any vector $u$, its $\ell_{2}$, $\ell_{1}$ and $\ell_{0}$ norms are defined respectively as $\|u\|_{2}=\sqrt{\sum_{i}u_{i}^{2}}$, $\|u\|_{1}=\sum_{i}|u_{i}|$, and $\|u\|_{0}=\sum_{i}\{u_{i}\neq 0\}$. Let $I_{m}$ denote the $m\times m$ identity matrix. For any matrix ${A=(a_{ij})\in\mathbb{R}^{n\times p}}$, $A(i,j)$ denote the $(i,j)$-entry of $A$, $A_{i\cdot}$ and $A_{\cdot j}$ denote the $i^{\text{th}}$ row and $j^{th}$ column of $A$ respectively. Throughout this paper, $\lambda_{i}(A)$ stands for the $i^{th}$ largest singular value of the matrix $A$ with ${\lambda_{\max}(A)=\lambda_{1}(A)}$, ${\lambda_{\min}(A)=\lambda_{p\wedge\text{rank}(A)}(A)}$. We also denote as $U_{K}(A)$ and $V_{K}(A)$ the left and right singular matrix from the rank-$K$ SVD of $A$. The frobenius, entry-wise $\ell_{1}$ norm and the operator norms of $A$ are denoted as $\left\|A\right\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, $\lVert A\rVert_{11}=\sum_{i,j}|a_{ij}|$, and ${\left\|A\right\|_{op}=\lambda_{1}(A)}$, respectively. The $\ell_{21}$ norm is denoted as $\big{\|}A\big{\|}_{21}=\sum_{i\in[n]}\|A_{i\cdot}\|_{2}$. For any positive semi-definite matrix $A$, $A^{1/2}$ denotes its principal square root that is positive semi-definite and satisfies $A^{1/2}A^{1/2}=A$. The trace inner product of two matrices $A,B\in\mathbb{R}^{n\times p}$ is denoted by $\langle A,B\rangle=\mathop{\sf Tr}(A^{\top}B)$. $A^{\dagger}$ denotes the pseudo-inverse of the matrix $A$ and $P_{A}$ denotes the projection matrix onto the subspace spanned by columns of $A$.
@@ 2.1 Assumptions
Let $\mathcal{G}=(\mathcal{D},\mathcal{E})$ denote an undirected graph induced by a known adjacency matrix on the $n$ documents in the corpus. The documents are represented as nodes $\mathcal{D}$, and $\mathcal{E}$ denotes the edge set of size $\lvert\mathcal{E}\rvert=m$. Throughout this paper, for simplicity, we will assume that $\mathcal{G}$ is binary, but our approach—as discussed in Section 2.3—can be in principle extended to weighted graphs. We denote the graph’s incidence matrix as $\Gamma\in\mathbb{R}^{m\times n}$ where, for the $d^{th}$ edge $e=(i,j),i<j$ between nodes $i$ and $j$ in the graph:
$\Gamma_{dg}=\begin{cases}1&\text{if }g=i\\ -1&\text{if }g=j\\ 0&\text{otherwise}\end{cases}$
It is easy to show that the Laplacian of the graph can be expressed in terms of the incidence matrix as $L=\Gamma^{\top}\Gamma$ (Hütter and Rigollet, 2016). Let $\Gamma^{\dagger}$ be the pseudo-inverse of $\Gamma$, and denote by $\mathbf{s}_{i},i=1\cdots m$ its columns, so that $\Gamma^{\dagger}=[\mathbf{s}_{1},\cdots,\mathbf{s}_{m}]$. We also define the inverse scaling factor of the incidence matrix $\Gamma$ (Hütter and Rigollet, 2016), a quantity necessary for assessing the performance of GpLSI:
@@
We focus on the estimation of the topic mixture matrix under the assumption that neighboring documents have similar mixture proportions: $W_{i\cdot}\approx W_{j\cdot}$ if $i\sim j$. This implies that the rows of $W$ are assumed to be smooth with respect to the graph $\mathcal{G}$. Noting that the difference of mixture proportions between neighboring nodes $i$ and $j$ ($e=(i,j)\in\mathcal{E}$) can be written as $(\Gamma W)_{l\cdot}=W_{i\cdot}-W_{j\cdot}$, this smoothness assumption effectively implies sparsity on the rows of the matrix $\Gamma W$.
Assumption 1 (Graph-Aligned mixture matrix)
The support (i.e, the number of non-zero rows) of the matrix $\Gamma W$ is small:
@@
The previous assumption is akin to assuming that the underlying mixture matrix $W$ is piecewise-continuous with respect to the graph $\mathcal{G}$, or more generally, can be well approximated by a piecewise-continuous function.
Our setting is not limited to connected document graphs. Denote $n_{\mathcal{C}}$ as the number of connected subgraphs of $\mathcal{G}$ and $n_{\mathcal{C}_{l}}$ as the number of nodes in the $l^{th}$ connected subgraph. Denote
@@
The error bound of our estimators will depend on the two quantities $n_{\mathcal{C}}$ and $n_{\mathcal{C}_{\min}}$. In the rest of this paper, we will assume that all connected components have roughly the same size: $n_{\mathcal{C}_{1}}\asymp\dots\asymp n_{\mathcal{C}_{n_{\mathcal{C}}}}$.
Assumption 2 (Anchor document)
For each topic $k=1,\dots,K,$ there exists at least one document $i$ (called an anchor document) such that $W_{ik}=1$ and $W_{il}=0$ for all $l\not=k$.
Remark 1
Assumption 2 is standard in the topic modeling community, as it is a sufficient condition for the identifiability of the topic and mixture matrices $A$ and $W$ (Donoho and Stodden, 2003). Beyond identifiability, we argue here that this could be interpreted as much as a definition as an assumption: Topics are interpretable only when associated with archetypes—that is, “extreme“ representations (in our case, anchor documents) — that illustrate the topic more expressively.
Assumption 3 (Equal Document Sizes)
@@
Assumption 4 (Condition number of $M$ and $W$)
There exist two constants $c>1$ and $c^{*}>0$ such that
@@
Assumption 5 (Assumption on the minimum word frequency)
We assume that the expected word frequencies $h_{j}$ defined as: $\forall j\in[p],h_{j}=\sum_{i=1}^{n}A_{ij}$ are bounded below by:
@@
Remark 2
Assumption 5 is a relatively strong assumption that essentially restricts the scope of this paper’s analysis to small vocabulary sizes (thereafter referred to as the “low-$p$” regime). Indeed, since $\sum_{j=1}^{p}\sum_{k=1}^{K}A_{jk}=K$, under assumption 5, it immediately follows that:
@@ 2.2 Estimation procedure: pLSI
Since the smoothness assumption (Assumption 1) pertains to the rows of the topic mixture matrix $W$, we build on the methodology introduced by Klopp et al. (2021). To the best of our knowledge, their approach achieves the best error bound to date on the matrix $W$. In this subsection, we provide a brief overview of their method.
When working in the oracle setting—that is, assuming we directly observe the true expected frequency matrix $M$ defined in Equation (2)—Klopp et al. (2021) propose a straightforward three-step procedure to recover the mixture matrix $W$. Specifically, let $U$ and $V$ be the left and right singular vectors obtained from the singular value decomposition of $M$, so that $M=U\Lambda V^{T}$. A critical insight from their work is that $U$ can be decomposed as
@@
where $H$ is a full-rank $K\times K$ matrix. From this decomposition, it follows that the rows of $U$—which can be viewed as $K$-dimensional embeddings of the documents—lie on the $K$-dimensional simplex $\Delta_{K-1}$. The simplex’s vertices, represented by the rows of $H$, correspond to the anchor documents (Assumption 2). Thus, identifying these vertices through any standard vertex-finding algorithm applied to the rows of $U$ will enable the estimation of $W$.
The procedure of Klopp et al. (2021) is summarized as follows:
@@
Compute the singular value decomposition of the matrix $M$, reduced to rank $K$, to obtain: $M=U\Lambda V^{T}$.
@@
Consider the rows of $U\in\mathbb{R}^{n\times K}$ as points in $\mathbb{R}^{K}$ on the $K$-dimensional simplex $\Delta_{K-1}$. To identify its vertices, Klopp et al. (2021) propose to use a vertex-hunting algorithm called the successive projection algorithm (SPA) (Araújo et al., 2001) on the rows of $U$. This algorithm returns the indices of the selected “anchor documents,” $J\subseteq[n]$ with $|J|=K$. Define $\widehat{H}=U_{J}$, where each row corresponds to one of the $K$ vertices of the simplex $\Delta_{K-1}$.
@@
$W$ can simply be recovered from $U$ and $\widehat{H}$ as:
$\widehat{W}={U}\widehat{H}^{-1}$ (7)
@@
Finally, $A$ can subsequently be estimated as: $\widehat{A}=\widehat{H}\Lambda V^{\top}.$
In the non-oracle case—that is, when we observe $X$ in Equation (2) instead of $M$—the procedure is adapted by plugging $X$ instead of $M$ in Step 1 and getting estimates of the singular vectors: $X=\widehat{U}\widehat{\Lambda}\widehat{V}^{\top}$. Klopp et al. (2021) prove the consistency of their estimator of $W$ under a similar set of assumptions as ours (Assumptions 2-4). The following corollary is significant because, to our knowledge, it offers one of the best error bounds on the estimation of the topic mixture matrix $W$ for pLSI.
Corollary 1 (Upper bound in the balanced case, from Klopp et al. (2021))
Let Assumptions 2-4 hold and let
$N$ $\geq CK^{5}\log(n+p)$
for some ${C}>0$ large enough. Then, with probability at least $1-o((n+p)^{-1})$, the matrix $\widehat{H}$ is non-degenerate and the output $\widehat{W}$ of the preconditioned SPOC algorithm (a version of the SPA algorithm) satisfies, for some constant ${C}_{0}>0$,
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq{\rm C_{0}}K\sqrt{\dfrac{n\log(n+p)}{N}},$
where $\mathcal{P}$ denotes the set of all permutation matrices.
However, this approach has two key limitations. First, the consistency of the estimator relies on having a sufficiently large number of words per document. The authors establish minimax error bounds, showing that the rate of any estimator’s error for $W$ is bounded below by a term on the order of $C\sqrt{\frac{n}{N}}$ (Theorem 3, Klopp et al. (2021)). In other words, the accurate estimation of $W$ requires that each document contains enough words. In many practical scenarios — such as the tumor microenvironment example mentioned earlier — this condition may not hold. However, we might still have access to additional information indicating that certain documents are more similar to one another.
Second, the method is relatively rigid and does not easily accommodate additional structural information, such as document-level similarities. Indeed, this method does not rely on convex optimization formulation to which we can add a regularization term, and the vertex-hunting algorithm does not readily incorporate metadata of the documents.

### v2

# Graph Topic Modeling for Documents with Spatial or Covariate Dependencies

@@ Abstract
Keywords: graph cross validation, graph regularization, latent dirichlet allocation, total variation penalty
@@ 1 Introduction
Originally developed to reduce document representations to low-dimensional latent semantic spaces, topic modeling has been successfully deployed for the analysis of count data in a number of applications, ranging from image processing (Tu et al. 2016, Zheng et al. 2015) and image annotation(Feng & Lapata 2010, Shao et al. 2009), to biochemistry (Reder et al. 2021), genetics (Dey et al. 2017, Kho et al. 2017, Liu et al. 2016, Yang et al. 2019), and microbiome studies (Sankaran & Holmes 2019, Symul et al. 2023).
@@
    Analyzing tumor microenvironments: In this context, slices of tumor samples are partitioned into smaller regions known as tumor microenvironments, where the frequency of specific immune cell types is recorded (Chen et al. 2020). Here, documents correspond to microenvironments, and words to cell types. The objective is to identify communities of co-abundant cells (topics), taken here as proxies for tumor-immune cell interactions and potential predictors of patient outcomes. In this setting, we assume that neighboring microenvironments share similar topic proportions. Since these microenvironments are inherently small, leveraging the spatial smoothness of the mixture matrix $W$ can significantly improve inference (Chen et al. 2020). We develop this example in further detail in Section 4.1.
@@
    Microbiome studies: Topic models have also been proven to be extremely useful in microbiome analysis (Sankaran & Holmes 2019, Symul et al. 2023). In this setting, the data consists of a microbiome count matrix recording the amount of different types of bacteria found in each sample. In this case, microbiome samples are identified to documents, with bacteria playing the roles of the vocabulary, and the goal becomes to identify communities of co-abundant bacteria (Sankaran & Holmes 2019). When additional covariate information is available (such as age, gender, and other demographic details), we can expect similar samples (documents) to exhibit similar community compositions (topic proportions).
@@
    The analysis of short documents, such as a collection of scientific abstracts or recipes: In this case, while recipes might be short, information on the origin of the recipe can help determine the topics and mixture matrix more accurately by leveraging the assumption that recipes of neighboring countries will typically share similar topic proportions. We elaborate on this example in greater detail in Section 4.3.
Prior works. Previous attempts to incorporate metadata in topic estimation have focused on the Bayesian extensions of latent Dirichlet allocation (LDA) model of Blei et al. (2001). By and large, these methods typically incorporate the metadata—often in the form of a covariate matrix—within a prior distribution (Blei & Lafferty 2006a, b, Roberts et al. 2014, Mcauliffe & Blei 2007). However, these models (a) are difficult to adapt to different types of covariates or information encoded as a graph, and (b) typically lack theoretical guarantees. Recent work by Chen et al. (2020) proposes extending LDA to analyze documents with known similarities by smoothing the topic proportion hyperparameters along the edges of the graph. However, this method does not empirically yield spatially smooth structures (see Sections  3.4 and  4), and significantly increases the algorithm’s running time.
In the frequentist realm, probabilistic latent semantic indexing (pLSI) has gained renewed interest over the past five years. Similar to LDA, it effectively models documents as bags of words but differs by treating matrices $A$ and $W$ as fixed parameters. In particular, new work by Ke & Wang (2017) and Klopp et al. (2021) suggest procedures to reliably estimate the topic matrix $A$ and mixture matrix $W$, characterizing consistency and efficiency through high-probability error bounds Although recent work has begun investigating the use of structures in pLSI-based topic models, most approaches have limited this to considering various versions of sparsity (Bing et al. 2020, Wu et al. 2023) or weak sparsity (Tran et al. 2023) for the topic matrix $A$. To the best of our knowledge, no pLSI approach has yet been proposed that effectively leverages similarities between documents nor characterizes the consistency of these estimators.
@@
    We propose a scalable algorithm based on a graph-aligned singular decomposition of the empirical frequency matrix $X$ to provide estimates of $W$ and $A$ (Section 2). Additionally, we develop a new cross-validation procedure for our graph-based penalty that allows us to choose the optimal regularization parameter adaptively (Section A of the Appendix).
@@
    We prove the benefits of the graph alignment procedure by deriving high probability upper bounds for both $W$ and $A$ in Section 3, which we verify through extensive simulations in Section 3.4.
@@
    Finally, we showcase the advantage of our method over LDA-based methods and non-structured pLSI techniques on three real-world datasets: two spatial transcriptomics examples and a recipe dataset in Section 4.
@@
Throughout this paper, we use the following notations. For any $t\in\mathbb{Z}_{+}$, $[t]$ denotes the set $\{1,2,...,t\}$. For any $a,b\in\mathbb{R}$, we write $a\vee b=\max(a,b)$ and ${a\wedge b=\min(a,b)}$. We use $\mathbf{1}_{d}\in\mathbb{R}^{d}$ to denote the vector with all entries equal to 1 and $\mathbf{e}_{k}\in\mathbb{R}^{d}$ to denote the vector with $k^{th}$ element set to 1 and 0 otherwise. For any vector $u$, its $\ell_{2}$, $\ell_{1}$ and $\ell_{0}$ norms are defined respectively as $\|u\|_{2}=\sqrt{\sum_{i}u_{i}^{2}}$, $\|u\|_{1}=\sum_{i}|u_{i}|$, and $\|u\|_{0}=\sum_{i}\mathbf{1}\{u_{i}\neq 0\}$. Let $I_{m}$ denote the $m\times m$ identity matrix. For any matrix ${A=(a_{ij})\in\mathbb{R}^{n\times p}}$, $A(i,j)$ denote the $(i,j)$-entry of $A$, $A_{i\cdot}$ and $A_{\cdot j}$ denote the $i^{\text{th}}$ row and $j^{th}$ column of $A$ respectively. Throughout this paper, $\lambda_{i}(A)$ stands for the $i^{th}$ largest singular value of the matrix $A$ with ${\lambda_{\max}(A)=\lambda_{1}(A)}$, ${\lambda_{\min}(A)=\lambda_{p\wedge\text{rank}(A)}(A)}$. We also denote as $U_{K}(A)$ and $V_{K}(A)$ the left and right singular matrix from the rank-$K$ SVD of $A$. The Frobenius, entry-wise $\ell_{1}$ norm and the operator norms of $A$ are denoted as $\left\|A\right\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, $\lVert A\rVert_{11}=\sum_{i,j}|a_{ij}|$, and ${\left\|A\right\|_{op}=\lambda_{1}(A)}$, respectively. The $\ell_{21}$ norm is denoted as $\big{\|}A\big{\|}_{21}=\sum_{i}\|A_{i\cdot}\|_{2}$. For any positive semi-definite matrix $A$, $A^{1/2}$ denotes its principal square root that is positive semi-definite and satisfies $A^{1/2}A^{1/2}=A$. The trace inner product of two matrices $A,B\in\mathbb{R}^{n\times p}$ is denoted by $\langle A,B\rangle=\mathop{\sf Tr}(A^{\top}B)$. $A^{\dagger}$ denotes the pseudo-inverse of the matrix $A$ and $P_{A}$ denotes the projection matrix onto the subspace spanned by columns of $A$.
@@ 2.1 Assumptions
Let $\mathcal{G}=(\mathcal{D},\mathcal{E})$ denote an undirected graph induced by a known adjacency matrix on the $n$ documents in the corpus. The documents are represented as nodes $\mathcal{D}$, and $\mathcal{E}$ denotes the edge set of size $\lvert\mathcal{E}\rvert=m$. Throughout this paper, for simplicity, we will assume that $\mathcal{G}$ is binary, but our approach—as discussed in Section 2.3—can be in principle extended to weighted graphs. We denote the graph’s incidence matrix as $\Gamma\in\mathbb{R}^{m\times n}$ where, for any edge $e=(i,j),i<j$ between nodes $i$ and $j$ in the graph, $\Gamma_{ei}=1$, $\Gamma_{ej}=-1$ and $\Gamma_{ek}=0$ for any $k\neq i,j.$ It is easy to show that the Laplacian of the graph can be expressed in terms of the incidence matrix as $L=\Gamma^{\top}\Gamma$ (Hütter & Rigollet 2016). Let $\Gamma^{\dagger}$ be the pseudo-inverse of $\Gamma$, and denote by $\mathbf{s}_{i},i=1\cdots m$ its columns, so that $\Gamma^{\dagger}=[\mathbf{s}_{1},\cdots,\mathbf{s}_{m}]$. We also define the inverse scaling factor of the incidence matrix $\Gamma$ (Hütter & Rigollet 2016), a quantity necessary for assessing the performance of GpLSI:
@@
We focus on the estimation of the topic mixture matrix under the assumption that neighboring documents have similar topic mixture proportions: $W_{i\cdot}\approx W_{j\cdot}$ if $i\sim j$. This implies that the rows of $W$ are assumed to be smooth with respect to the graph $\mathcal{G}$. Noting that the difference of mixture proportions between neighboring nodes $i$ and $j$ ($e=(i,j)\in\mathcal{E}$) can be written as $(\Gamma W)_{d\cdot}=W_{i\cdot}-W_{j\cdot}$, this smoothness assumption effectively implies sparsity on the rows of the matrix $\Gamma W$.
Assumption 1 (Graph-Aligned mixture matrix).
The support (i.e, the number of non-zero rows) of the difference matrix $\Gamma W=\left(W_{i\cdot}-W_{j\cdot}\right)_{(i,j)\in\mathcal{E}}$ is small:
@@
The previous assumption is akin to assuming that the underlying mixture matrix $W$ is piecewise-continuous with respect to the graph $\mathcal{G}$, or more generally, that it can be well approximated by a piecewise-continuous function.
Our setting is not limited to connected document graphs. Denote $n_{\mathcal{C}}$ the number of connected subgraphs of $\mathcal{G}$ and $n_{\mathcal{C}_{l}}$ the number of nodes in the $l^{th}$ connected subgraph. Let $n_{\mathcal{C}_{\min}}$ be the size of the smallest connected component:
@@
The error bound of our estimators will depend on both $n_{\mathcal{C}}$ and $n_{\mathcal{C}_{\min}}$. In the rest of this paper, we will assume that all connected components have roughly the same size: $n_{\mathcal{C}_{1}}\asymp\dots\asymp n_{\mathcal{C}_{n_{\mathcal{C}}}}$.
Assumption 2 (Anchor document).
For each topic $k=1,\dots,K,$ there exists at least one document $i$ (called an anchor document) such that $W_{ik}=1$ and $W_{ik^{{}^{\prime}}}=0$ for all $k^{{}^{\prime}}\not=k$.
Remark 1.
Assumption 2 is standard in the topic modeling community, as it is a sufficient condition for the identifiability of the topic and mixture matrices $A$ and $W$ (Donoho & Stodden 2003). Beyond identifiability, we contend that the “anchor document” assumption functions not only as a sufficient condition for identifiability but also as a necessary condition for interpretability: Topics are interpretable only when associated with archetypes—that is, “extreme“ representations (in our case, anchor documents)—that illustrate the topic more expressively.
Assumption 3 (Equal Document Sizes).
@@
Assumption 4 (Condition number of $M$ and $W$).
There exist two constants $c,c^{*}>0$ such that
@@
Assumption 5 (Assumption on the minimum word frequency).
We assume that the expected word frequencies $h_{j}$ defined as: $\forall j\in[p],h_{j}=\sum_{k=1}^{K}A_{kj}$ are bounded below by:
@@
Remark 2.
Assumption 5 is a relatively strong assumption that essentially restricts the scope of this paper’s analysis to small vocabulary sizes (thereafter referred to as the “low-$p$” regime). Indeed, since $\sum_{j=1}^{p}\sum_{k=1}^{K}A_{kj}=K$, under Assumption 5, it immediately follows that:
@@ 2.2 Estimation procedure: pLSI
Since the smoothness assumption (Assumption 1) pertains to the rows of the document-topic mixture matrix $W$, we build on the pLSI algorithm developed by Klopp et al. (2021). In this subsection, we provide a brief overview of their method.
When we assume we directly observe the true expected frequency matrix $M$ defined in Equation (2), Klopp et al. (2021) propose a fast and simple procedure to recover the mixture matrix $W$. Specifically, let $U\in\mathbb{R}^{n\times K}$ and $V\in\mathbb{R}^{p\times K}$ be the left and right singular vectors obtained from the singular value decomposition (SVD) of the true signal $M\in\mathbb{R}^{n\times p}$, so that $M=U\Lambda V^{T}$. A critical insight from their work is that $U$ can be decomposed as:
@@
where $H$ is a full-rank, $K\times K$-dimensional matrix. From this decomposition, it follows that the rows of $U$, which can be viewed as $K$-dimensional embeddings of the documents, lie on the $K$-dimensional simplex $\Delta_{K-1}$. The simplex’s vertices, represented by the rows of $H$, correspond to the anchor documents (Assumption 2). Thus, identifying these vertices through any standard vertex-finding algorithm applied to the rows of $U$ will enable the estimation of $W$. The procedure of Klopp et al. (2021) can be summarized as follows:
@@
Compute the singular value decomposition (SVD) of the matrix $M$, reduced to rank $K$, to obtain: $M=U\Lambda V^{T}$.
@@
Apply the successive projection algorithm (SPA) (Araújo et al. 2001) (a vertex-hunting algorithm) on the rows of $U$. This algorithm returns the indices of the selected “anchor documents,” $J\subseteq[n]$ with $|J|=K$. Define $\widehat{H}=U_{J\cdot}$, where each row corresponds to one of the $K$ vertices of the simplex $\Delta_{K-1}$.
@@
$W$ can simply be recovered from $U$ and $\widehat{H}$ as
$\widehat{W}={U}\widehat{H}^{-1}.$ (7)
@@
Finally, $A$ can subsequently be estimated as $\widehat{A}=\widehat{H}\Lambda V^{\top}.$
In the noisy setting, the procedure is adapted by plugging the observed frequency $X$ instead of $M$ in Step 1 and getting estimates of the singular vectors: $X=\widehat{U}\widehat{\Lambda}\widehat{V}^{\top}$. Under a similar set of assumptions as ours (Assumptions 2-4), Theorem 1 of (Klopp et al. 2021) states that the error of $\widehat{W}$ is such that: $\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq{\rm C_{0}}K\sqrt{n\log(n+p)/N}$, where $\mathcal{P}$ denotes the set of all permutation matrices. Their analysis provides one of the best error bounds on the estimation of the topic mixture matrix $W$ for pLSI.
However, this approach has two key limitations. First, the consistency of their estimator relies on having a sufficiently large number of words per document, $N$. In particular, a necessary condition for the aforementioned results to hold is that $N\geq CK^{5}\log(n+p)$. The authors establish minimax error bounds, showing that the rate of any estimator’s error for $W$ is bounded below by a term on the order of $O(\sqrt{n/N})$ (Theorem 3, Klopp et al. (2021)). In other words, the accurate estimation of $W$ requires that each document contains enough words. In many practical scenarios—such as the tumor microenvironment example mentioned earlier—this condition may not hold. However, we might still have access to additional information indicating that certain documents are more similar to one another.
Second, the method is relatively rigid and does not easily accommodate additional structural information, such as document-level similarities. Indeed, this method does not rely on a convex optimization formulation to which we could simply add a regularization term, and the vertex-hunting algorithm does not readily incorporate metadata of the documents.